It is 1100 miles from Hacienda Heights to Seattle, Washington. The plane that Cerena takes flies 250 mph. If Cerena leaves at 7:30 a.m, when does she arrive in Seattle(to the nearest minute)
Answer by oberobic(2304)

d = r*t is the basic distance equation
.
1100 miles = 250 mph * t
.
t = 1100/250 = 22/5 hr
.
22/5 hr = 4 2/5 hr
.
2/5 of an hr = 2/5 hr * 60 min/hr = 24 min
.
t = 4 hr 24 min
.
7:30 a.m. + 4:24 = 11:54 a.m. is the time of arrival.
.
Done.